# An “opinion reproduction number” for infodemics in a bounded-confidence content-spreading process on networks

### v1

@@ VII Conclusions and discussion
We examined “infodemics”, in the form of large content-spreading cascades, in a bounded-confidence content-spread model on configuration-model networks. To do this, we defined an “opinion reproduction number” that is analogous to the basic reproduction number from disease dynamics. By examining whether opinion spread is below or above the opinion reproduction number — such that it is below or above a corresponding “infodemic threshold” — we investigated when opinions die off and when they propagate widely in a population of agents. We quantified the size and structure of content spread by measuring several properties of dissemination trees — tree size, tree width, the length of the longest adoption paths, and structural virality — and we thereby illustrated how network structure and spreading-model parameters affect how content spreads. We found that larger networks, larger expected mean degrees, larger receptiveness, and content states near the mean of opinion distributions all promote larger total numbers of content shares, including the possibility of sharing the content to many new nodes in a single time step. Additionally, when the expected mean degree and receptiveness levels are small but place the spreading process above the infodemic threshold, there can be longer dissemination trees than when the expected mean degree or receptiveness are larger, even though the total number of content shares is smaller. This indicates that content that is spreading farther but less widely from a source agent (i.e., the agent that originally posts a piece of content).
@@
There are many interesting ways to extend our work. A particularly relevant one is to examine the effects of purposeful choices of source nodes, such as to try to promote influence maximization Peng et al. (2018). One can also examine the effects of competing social contagions Gleeson et al. (2014), which perhaps are spread by agents with different political perspectives or agendas. One can also incorporate content mutation Adamic et al. (2016) (including through opinion averaging, as in classical BCMs such as the Deffuant–Weisbuch model Deffuant et al. (2000)) and examine how that changes infodemic thresholds and spreading patterns. Another worthwhile future direction is to adapt the recently developed “distributed reproduction numbers” She, Paré, and Hale (2023) from disease spreading to content spreading. It is also desirable to extend our study of content spreading to other network structures. One key direction is to consider more complicated network models (such as generalizations of configuration models that incorporate various types of heterogeneities Melnik et al. (2014)) and real-world networks. It is also natural to extend our investigation to more complicated types of networks, such as multilayer networks (which allow multiple types of social connections and communication channels) Kivelä et al. (2014), hypergraphs (which allow simultaneous interactions between three or more agents) Battiston et al. (2020), and adaptive networks (e.g., to incorporate changing relationship structures such as “unfollowing” or “unfriending” on social-media platforms) Berner et al. (2023).
@@ Acknowledgements.
We thank Jonas Juul for helpful discussions. HZB was funded in part by the National Science Foundation (grant number DMS-2109239) through their program on Applied Mathematics. MAP acknowledges financial support from the National Science Foundation (grant number 1922952) through their program on Algorithms for Threat Detection (ATD). MAP thanks David Campbell for his friendship and many discussions over the years.
VIII A few words about David Campbell (by Mason A. Porter)
@@
I first interacted with David when I was a postdoc at Georgia Tech. I was naively trying to publish a modified version of my doctoral thesis Porter (2002) as a review article, and I submitted it to the journal Physics Reports. David, who I had yet to meet in person, was the handling editor of my submission. My paper was rejected, but what stood out in that process was the carefulness and decency in which David handled the process. All I want with a manuscript submission — and all anyone should ever want with a manuscript submission, as both a necessary and sufficient condition — is to be treated fairly and with dignity. David saw a teaching opportunity, and he could see that I had things to learn, and he went beyond the call of duty and even asked the referee to disclose his identity to help with that teaching. By contrast, as we all see repeatedly, many journal editors seem to be bean counters, rather than actual editors. It wasn’t my ideal outcome, but it was the correct outcome, and I learned a lot from it. In my own editing roles, I have tried to draw on the approach and lessons that David exemplified in this story.
@@
David continued looking out for me, as he has looked out for many others. Having — I think? — still not yet met in person, David invited me to give a talk as part of a special session at the 2005 American Physical Society (APS) March Meeting about the celebrated Fermi–Pasta–Ulam–Tsingou (FPUT) problem and contribute an article about the dynamics of Bose–Einstein condensates Porter et al. (2005) to a special issue in Chaos about the FPUT problem. David’s talk had a pop-music reference (which was about the Eagles, if memory serves), and he mentioned that he figured that I’d appreciate the music reference. (He was right.) Four of us in the special session eventually coauthored a popular article about the FPUT problem Porter, Zabusky, and Campbell (2009). I eventually wrote a variety of papers on the FPUT problem (and on other related problems), including one Nelson, Porter, and Choubey (2018) that turned out to be refereed by David and one of his students. It was a particularly fair and thorough referee report — of the type that scientists dream of getting that indicates that sometimes the reviewing process actually works the way that it it supposed to — that my coauthors were spending time trying to address. I remember telling David by e-mail during our revision process to tell him about our paper because I figured that he would be interested in this paper, which concerns heterogeneities in FPUT lattices. My e-mail to David either was about something else entirely or was because I had just seen a paper of his on a related topic. In any event, without any idea that David was the referee, I told him in my e-mail about our paper and that we were currently dealing with a tough-but-fair referee report. I found out from David’s response that he and his student had written that report. Once again, David was doing things in a way that helps authors while upholding rigorous standards. In other words, David once again did things in the way that they are supposed to be done
@@
A recurring theme in my stories is David looking out for a junior scientist, fair treatment while retaining rigor, and doing things the right way in scientific editing and review. We all have reviewing horror stories and publishing can cause many frustrations. But some scientists consistently do things the right way, and David is one of them.
@@
The best complement that somebody like me who comes from Jewish heritage can pay to another person is to call them a “mensch” (which, essentially, is calling somebody a decent human being, as they are a person of integrity and honor). David Campbell is a mensch.

### v2

@@ VII Conclusions and discussion
We examined “infodemics”, in the form of large content-spreading cascades, in a bounded-confidence content-spreading model on configuration-model networks. To do this, we defined an “opinion reproduction number” that is analogous to the basic reproduction number of disease dynamics. By examining whether the opinion of the content and the receptiveness of individuals yield an opinion reproduction number that is larger than $1$, we investigated when content propagates widely in a network of agents and when it does not.
@@
We quantified the size and structure of content spread by measuring several properties of dissemination trees — tree size, tree width, the length of the longest adoption paths, and structural virality — and we thereby illustrated how network structure and spreading-model parameters affect how content spreads. We found that larger networks, larger expected mean degrees, larger receptiveness, and content states near the mean of opinion distributions all promote larger total numbers of content shares, including the possibility of sharing the content to many new nodes in a single time step. Additionally, when the expected mean degree and receptiveness levels are small but place the spreading process above the infodemic threshold, there can be longer dissemination trees than when the expected mean degree or receptiveness are larger, even though the total number of content shares is smaller. This indicates that content is spreading farther but less widely from a source agent (i.e., an agent that originally posts a piece of content).
@@
There are many interesting ways to extend our work. A particularly relevant extension is to examine the effects of purposeful choices of source nodes, such as to try to promote influence maximization Peng et al. (2018); Kempe, Kleinberg, and Tardos (2003). One can also examine the effects of competing social contagions Gleeson et al. (2014), which perhaps are spread by agents with different political perspectives or agendas. Another worthwhile future direction is to adapt the recently developed “distributed reproduction numbers” She, Paré, and Hale (2023) from disease spread to content spread. It is also desirable to extend our study of content-spreading dynamics to other network structures. One key direction is to consider more complicated network models (such as generalizations of configuration models that incorporate various types of heterogeneities Melnik et al. (2014)) and real-world networks. It is also natural to extend our investigation to more complicated types of networks, such as multilayer networks (which allow multiple types of social connections and communication channels) Kivelä et al. (2014), hypergraphs (which allow simultaneous interactions between three or more agents) Battiston et al. (2020), and adaptive networks (e.g., to incorporate changing relationship structures such as “unfollowing” or “unfriending” on social-media platforms) Berner et al. (2023).
Other viable extensions of our model directly exploit the link that it establishes between opinion dynamics and percolation processes. It seems particularly exciting to study scenarios with rich interplays between the opinion of content and the opinions of agents. For example, one can incorporate mutations of content opinions Friggeri et al. (2014); Adamic et al. (2016) through averaging the the opinion of content with the opinions of the agents that spread it. A related extension entails allowing agent opinions to evolve with time through their interactions with each other (as in classical BCMs, such as the Deffuant–Weisbuch model Deffuant et al. (2000)) while content with a fixed opinion spreads on a network. It will be fascinating to explore how these extensions affect the properties of infodemic thresholds and spreading patterns.
@@ Acknowledgements.
We thank Abhinav Chand, Jonas Juul, and two anonymous referees for helpful comments. HZB was funded in part by the National Science Foundation (grant number DMS-2109239) through their program on Applied Mathematics. MAP acknowledges financial support from the National Science Foundation (grant number 1922952) through their program on Algorithms for Threat Detection (ATD). MAP thanks David Campbell for his friendship and many discussions over the years.
Appendix A A few words about David Campbell (by Mason A. Porter)
@@
I first interacted with David when I was a postdoc at Georgia Tech. I was naively trying to publish a modified version of my doctoral thesis Porter (2002) as a review article, and I submitted it to the journal Physics Reports. David, who I had not yet met in person, was the handling editor of my submission. My paper was rejected, but what stood out in that process was the carefulness and decency in which David handled the process. All that I want with a manuscript submission — and all that anyone should ever want with a manuscript submission, as both a necessary and sufficient condition — is to be treated fairly and with dignity. David saw a teaching opportunity, and he could see that I had things to learn; he went beyond the call of duty and even asked the referee to disclose his identity to help with that teaching. By contrast, as we all see repeatedly, many journal editors seem to be bean counters, rather than actual editors. The rejection of my paper wasn’t my ideal outcome, but it was the correct outcome, and I learned a lot from it. In my own editing roles, I have tried to draw on the approach and lessons that David exemplified in this story.
@@
David continued looking out for me, as he has looked out for many others. Having — I think? — still not yet met in person, David invited me to give a talk as part of a special session at the 2005 American Physical Society (APS) March Meeting about the celebrated Fermi–Pasta–Ulam–Tsingou (FPUT) problem and to contribute an article about the dynamics of Bose–Einstein condensates Porter et al. (2005) to a special issue in Chaos about the FPUT problem. David’s talk had a pop-music reference (which was about the Eagles, if memory serves), and he mentioned that he figured that I’d appreciate the music reference. (He was right.) Four of us in the special session eventually coauthored a popular article about the FPUT problem Porter, Zabusky, and Campbell (2009). I eventually wrote a variety of papers on the FPUT problem (and on related problems), including one Nelson, Porter, and Choubey (2018) that turned out to be refereed by David and one of his students. It was a particularly fair and thorough referee report — of the type that scientists dream of getting and that indicates that sometimes the reviewing process actually works the way that it it supposed to — that my coauthors were spending time trying to address. I remember e-mailing David during our revision process to tell him about our paper, which concerns heterogeneities in FPUT lattices, because I figured that he would be interested it. My e-mail to David either was about something else entirely or was because I had just seen a paper of his on a related topic. In any event, without having any notion that David was our referee, I told him in my e-mail about our paper and that we were currently dealing with a tough-but-fair referee report. I found out from David’s response that he and his student had written that report. Once again, David was doing things in a way that helps authors while upholding rigorous standards. In other words, David once again did things in the way that they are supposed to be done.
@@
A recurring theme in my stories is David looking out for a junior scientist, treating people fairly while simultaneously ensuring rigorous standards, and doing things the right way in scientific editing and reviewing. We all have reviewing horror stories, and publishing can cause many frustrations. However, some scientists consistently do things the right way, and David is one of them.
@@
The best complement that somebody like me who comes from Jewish heritage can pay to another person is to call them a “mensch” (which, essentially, is saying that somebody is a decent human being, as they are a person of integrity and honor). David Campbell is a mensch.